OpenStudy (caozeyuan):

we know that pI is the pH at which amino acid or other molecule exist at zwitterionic form. so the fraction of zwiitterion at pI is 1. Does this imply that given pH=1/2pI, zwitterion fraction is 1/2

OpenStudy (caozeyuan):

@Photon336 any idea?

OpenStudy (photon336):

Well, let's start off like this: the PI is the average of the two pKa for the amino and carboxylate groups. \[pI = 0.5(pKa_{1} + pKa_{2})\] To go back to general chemistry, let's look at what happens when the pH = the pKa. This fancy looking equation below: \[pH = pKa + \log[\frac{ A^{-} }{ HA }]\] for the pH to equal the pKa the concentration of [A-] our deprotonated acid must equal the concentration of the deprotonated form [HA] \[pH = pKa + \log[1], \log[1] = 0 \] now for amino acids I would guess that at the PI the amino acid wont migrate in an electric field because there's no net charge on it.

OpenStudy (caozeyuan):

Why is there [math processing error] instead of equations?!

OpenStudy (photon336):

Let me just write it on paper. and post it as attachement

OpenStudy (caozeyuan):

great, thank you so much

OpenStudy (photon336):

I don't see how the pH would be 1/2 the PI. because I think the pI is the PH at which the zwitterion would form. if i'm also correct, a fraction of half the zwitter ion would mean a fraction that's uncharged. I tried to make a pdf based on my understanding of it. because at the pI that's the pH where the zwitter ion would predominate and it's fraction should be 1 I guess that makes sense to me.